OpenStudy (anonymous):

Two urns contain white balls and yellow balls. The first urn contains 8 white balls and 8 yellow balls and the second urn contains 10 white balls and 3 yellow balls. A ball is drawn at random from each urn. What is the probability that both balls are white?

OpenStudy (anonymous):

@mathmale @e.mccormick

OpenStudy (e.mccormick):

OK, so multiple, independent events. So, do you know what process you use to find multiple independent events? The word multiple is a really good clue.

OpenStudy (anonymous):

No I don't.

OpenStudy (e.mccormick):

Well, what math word is close to multiple. You learned it really early on!

OpenStudy (anonymous):

Multiply?

OpenStudy (e.mccormick):

Yep! Find each independent event and multiply them.

OpenStudy (e.mccormick):

So if I had 2 decks of cards and I wanted to see if I would get a king out of the first and a heart out of the second it would be: \(\dfrac{4}{52}\cdot \dfrac{13}{52}\)

OpenStudy (anonymous):

Oh that's 52/2704.

OpenStudy (anonymous):

I had to refresh the page

OpenStudy (e.mccormick):

You need to simplify: \(\dfrac{4}{52}\cdot \dfrac{13}{52}=\dfrac{1}{13}\cdot \dfrac{1}{4}=\dfrac{1}{52}\) Yes, I chose single card and suit because it comes out nice like that. Do te same sort of thing with your jars.

OpenStudy (anonymous):

Ok so I have 1/2 *10/13 = 10/26. Is this correct?

OpenStudy (e.mccormick):

Urns. Whatever. So individual probabilities get multiplied when doing multiple, independent events. Or for short: Multiple events means multiply the events! Simplify yours.

OpenStudy (anonymous):

Ok so 5/13?

OpenStudy (e.mccormick):

Yep.

OpenStudy (anonymous):

Thank you :)